# Model Lakes

### v1

@@ 4. Model Lakes: Vision
Related Model Search. Another important problem is related model search. A model is considered similar or related to another if they have some degree of commonality in one or more combinations of the following categories: datasets, training algorithms, behavior, or purpose. One approach to addressing related-model search was explored by Lu et al. (2023), which searches for image generative models by using the behavior of another image model as a query. We propose extending such model-query searches to all models in a model lake to help users identify related models when exploring the model in question. This can then be used as an informative analytics visualization as shown in Figure 1.
Documentation Verification and Auditing. It is crucial to be able to verify that manually created documentation is correct and complete. When documentation is missing or incomplete, can we learn it? This is similar to discovering and annotating metadata in data lakes (Sawadogo and Darmont, 2021; Zhao et al., 2021; Korini and Bizer, 2023; Fan et al., 2023a). If there is a documentation, the goal would be to verify them and add any information that was left out. Currently, Mithril AICert (Security, [n.d.]a), for instance, runs an initiative to provide certification to indicate that a model is trained with the specified algorithm and with the specified training dataset. However, there are a couple of caveats. First, this initiative is still under development and hence, cannot be used for production at the time of writing. Second, this mechanism is dependent on creators signing up to gain such certification. Furthermore, based on the limitations provided on AICert (Security, [n.d.]c) website, the training code and data itself cannot be audited for threats such as “backdoors injected into a model by the code or poisonous data.” Lastly, it also states that AICert also needs to be inspected since we are currently trusting that it is running at a safe server and does not have a malignant app from which users have to run their code on the training data. Hence, there are a couple of broad research questions to explore within the document-centric verification realm – What information do we need for a model to be reproducible? If the information is provided, how can we verify the information without being compromised in the process? If the information is not provided, how can we gather the necessary information from the models themselves? How can we automate this process? While developing models, there are systems like MISTIQUE (Vartak et al., 2018), that allows developers to perform model diagnosis, i.e., the process of examining machine learning model performance, and aiming to understand where the model excels and where it falls short. MISTIQUE can do so by capturing and storing the model intermediates, which are generally data artifacts that relate to the model including input data, prediction values, and data representations produced by model. These intermediate values can then be queried by the developers to diagnose the model to understand its performance. However, such frameworks are limited to white-box models, which are models where users can have full access including access to weights, activation, and gradients. However, we also need to consider documentation verification for other variations of model access such as black-box access, where users can only see the model’s output for their queries. Currently, there is no rigorous scheme present, especially for AI auditing (Casper et al., 2024).
Data Citation. Data Citation helps stakeholders identify the source, ownership, and authorship of the data used for a particular analysis. It is important to use proper data citation because the structure and contents of database can evolve (Buneman et al., 2016). Hence, this problem is also extended to machine learning research, since a big part of model training is its dataset. Based on a study conducted on 13K model cards (Liang et al., 2024), researchers found that if a training section of the model card is filled out, they generally include the name of the dataset, volume and characteristics of the dataset. This is not sufficient since the exact version of the dataset is usually not specified and the download links to the datasets may become stale. Furthermore, without any standardization and automated process on how to reference these datasets, many model cards might not contain such information or could feature varied statements regarding their referencing, as exemplified in the case shared by (Buneman et al., 2016).
In addition to citing datasets used during the model’s training and evaluation stages, we also need to consider how to manage data generated by these models. It is important to do so because users can use the generated information to train another model. One of the solutions for at least recognizing a generated data output is use watermarks (Kirchenbauer et al., 2023). Another scenario is when data is generated by a model and is being used by a consumer. It is important to know the model’s version and input prompts to the model to improve the chances of recreating this dataset. Hence, there is a need to advance data citation research for models to build an ecosystem where crucial information about related databases are recorded in both data and model lakes.
Provenance. A general research question for model provenance can be as follows: Given a potential source model, , and a single target model, , is the source model of ? If so, how is transformed into ? These transformations could be related to dataset being updated, training algorithms being changed, and/or changes in requirements of the model indicated by the developer as described by Mei et al. (2022). Inspired from data provenance, we can also explore where-, how- and why-provenance with respect to models. To get a comprehensive understanding of model provenance, it is important to know what data was used to train it. Using data citation helps us collect this information effectively. There are, however, alternative methods (Liu et al., 2023) that rely on intermediate model training checkpoints and not training data to at least verify model ownership. In terms of exploring why-provenance, we can try to extract datasets used or changed with the source and target models. For instance, there are a range of facts that can be edited using model editing techniques. A 1-fact edited model would mean that there was one fact used whose association is different from its source model. For instance, if ROME algorithm (Meng et al., 2022) was used to change the location of the Eiffel Tower from Paris to Rome, we intend to edit one sentence representation from “Eiffel Tower is located in Paris” to “Eiffel Tower is located in Rome”. Hence, the resulting model would be a 1-fact edited model of its source model. The first comparison that can be made within this sub-area, i.e., can we determine whether one model is 1-fact edited model of another. A general approach to understanding such models would be to reverse engineer the model’s prediction of outputs, which can lead to circuit discovery (Wang et al., 2022; Conmy et al., 2023; Elhage et al., 2021) or even understanding how to unlearn knowledge learned in models (Jang et al., 2023; Eldan and Russinovich, 2023; Gandikota et al., 2023; Belrose et al., 2023; Kumari et al., 2023). We can extend the idea of such reverse engineering techniques to understand the process of 1-fact editing algorithms such as ROME on such models. Once such an ”edit or change” detector can be created, we could modify and extend its application. From model editing, we can then move on to model partial-tuning where a subset of the models’ parameters are used for training. We can look into the changed values for those parameters and understand what changed between its source and its current form. After building detectors for such cases, we can extend them further to understand the difference between one model that was fine-tuned from another model. Within fine-tuning, there are special cases of one-shot and few-shot training. Understanding these cases could help us build better detectors for the completely fine-tuned models where all model parameter values are usually changed. By creating such detectors, we can explain the transformation between a source and a target model, which is required for an exhaustive understanding of the target model’s provenance.
Model Version Search. While model provenance is a research challenge for identifying and explaining the transformation of one model, an extension to this challenge could be to identify versions given a set of models, and a set of datasets, . A model is considered a version of another when one can be transformed into another (such as through updates in dataset and/or training algorithm) and have the same true source model (i.e., the same first version). Apart from the approaches mentioned earlier, an alternative way to explore versioning is to understand how different entities have similar and/or different information present in various models. This could, for instance, be a post-hoc analysis through creation of knowledge graphs based on particular facts in data lakes that we know change from time to time, such as the President of the United States. By doing so, we could create a version history for data through models published at various time period, and by extension, identify the chronological order of the models based on their knowledge space.
Finally, by being able to pursue any direction indicated in the roadmap above, we can create a dataset or benchmark for such research so that the research community can build upon and create better methods for detecting such model differences. In addition to such benchmarks, it is also crucial to create metrics to understand how well these methods perform. By doing so, we can also increase outside-the-box access of AI models, where obtaining verified contextual information about the model can provide better transparency and evaluation for AI audits (Casper et al., 2024). Lastly, the solution to these methods would need visualization mechanisms to illustrate how these models and data are related. For instance, maybe we can have a hierarchical/network map of how different models are connected together? The next set of questions would be which parts of this mechanism can be automated and which ones would require a human-in-the-loop.
5. Granularity of Model Differences
There are various levels of granularity of model differences to consider. One aspect of granularity is explaining the transformation of a model. A low level granularity to identify the difference would be to identify the set of facts or data that were changed or were affected. A medium level of granularity could look like identifying the parameters that were changed and what dataset was it most likely trained on. A high level of granularity would be to just identify whether a model-in-question was fine-tuned from other model.
Another aspect of granularity is the level of information used to understand the model differences – did we use just the model weights? do we need their datasets? did we also use their documentation, i.e., their model cards? If we can find the difference between models by just using their model weights, then we can consider this difference to be a low-level information granularity. If we assume that we know at least one of the models’ datasets, then we can consider the difference to be a medium level. Lastly, if we depend on documentation to understand a pair of models’ differences, then we would consider it as a high-level of granularity. This can also scale the level of trust or confidence we can have for each model difference or similarity identification. For instance, identifying the difference between two models using just documentation has low level of confidence because we rely on trusting that the documentation is accurate. Furthermore, the accessibility of documentation can be limited because there is currently no structured way of documenting a model’s training process and inputs. However, if we rely on just the model’s weights, we can trust the identification more since it does not have an extra layer of information to process and understand the model.
Caveats and Considerations. There are a couple concerns with respect to extracting information about a model with low level granularity. If we were able to extract specific training data facts using just model weights, it can also be considered a privacy attack, which would mean that the model is unsafe in terms of privacy. Hence, in terms of providing granularity/verification scores as shown in Figure 1, it would have a high data verification score, but a low confidence in privacy. Another challenge with model lake tasks is differentiating between a predictive path and a deterministic path. For instance, in the case of identifying the source model of a certain model (model provenance task), we might find a couple of models that meet the possibility of it being a source model. In such cases, a lower than normal granularity score can be assigned unless a potential source model option is backed by another source of information, such as the model card.
6. MODEL LAKES: APPLICATION
Similar to a data lake, a model lake can be defined as a system or repository of many heterogeneous models and their related data stored in their natural format, for example, as files of model parameters. As the community progresses along the proposed directions, we can develop an intuitive navigation of a model lake, as depicted in Figure 1. This improvement can be helpful for addressing the Model Selection Problem shared in Example 1. In the left-side of Figure 1, we show that a user can search for models by typing out the task and type of data they are interested in working with. Based on this input, there can be an algorithm to semantically parse this text and find closely related task and data tags. Based on these tags, the search engine can share the most relevant models. Upon clicking a model, the model lake interface can lead to a model card page that may look similar to the right-side of the same figure. In this page, the users can view the model card description as well as what the model’s previous, next, source, and related models. In addition, they can also view what the training data was in order to understand if it was trained on desired text. If the training data is not specified, another way of understanding whether the model has a semantic understanding of the type of text the user cares about (for instance, legal jargon in the case of Example 1) is by viewing a semantic knowledge graph of the model and filtering for ‘legal’ within that graph. Within the model card, there can also be a privacy scale of low, medium, and high to gauge whether the model tends to share private information that was used in training data. Lastly, it could also have a set of verification scores (verifiable documentation task) to gauge how well this model is verified. The verification is categorized into data, code, and document verification. This scale would vary based on the level of granularity that they were verified with. 0 would indicate that they haven’t been verified and 1-5 would represent high level to low level verification where where high level implies less aspects of the model verified verses low level indicates that most, if not all, the verification was done using the model weight values.

### v2

@@ 4. State-of-the-art in Model Mgmt
Another important problem is related model search. One approach to addressing related model search was explored by Lu et al. (2023), who search for image generative models by using the behavior of another image model as a query. We propose extending such model as query searches to the model lake to help users identify related models when exploring the model in question.
Attribution. An important line of inquiry considers questions related to provenance or attribution (Mei et al., 2022; Mu et al., 2023). These issues pose similar questions to those studied in the database community (Buneman and Tan, 2018): from where did a generated fact derive or why was a predication made? Building on the concept of data provenance (Buneman et al., 2001), we can extend the notions of why-, where-, and how-provenance to models, under the name of model attribution. Similar research questions are also posed in model interpretability as part of local and global explanations. As in data provenance, the main issue is in whittling down all the provenance associated with some process (such as a full model) into simple, but useful explanations (Buneman and Tan, 2018). More than two decades of research in the data management community have produced elegant and simple, yet powerful, models for data provenance when the process is a query (Buneman et al., 2008, 2016) or workflow (Moreau et al., 2013; Moreau and Groth, 2013). In large model attribution, the goal cannot be to understand and track all inputs (data, hyperparameters, code, training regimes, …) that were used to create a specific model output. Hence, the challenge is to find meaningful sets of concepts that can be tracked efficiently and that provide important insights into model behavior.
The training data attribution problem is nontrivial because every aspect of training has the potential to impact every decision of a model. A variety of methods have been developed to estimate influence of training data on model behavior (Koh and Liang, 2017; Grosse et al., 2023; Wang et al., 2023). The current approaches require extensive use of training data as well as costly analysis of model intrinsics. Another approach to the problem is to apply techniques such as membership inference analyses, or membership inference attacks (Shi et al., 2024) that ask the question of whether a specific training data item is present in the training data , or training data reconstruction methods that extract sets of items from the training data (Carlini et al., 2021; Schwarzschild et al., 2024).
Insight about the attribution of model behavior can also be studied through model interpretability methods: extraction of relevant knowledge from a model concerning relationships either contained in data or learned by a model (Murdoch et al., 2019). Several works have surveyed techniques and research questions in this area (Mueller et al., 2024; Singh et al., 2024; Luo and Specia, 2024; Zhao et al., 2024; Räuker et al., 2022; Doshi-Velez and Kim, 2017; Gilpin et al., 2018). These approaches can be broadly categorized into the following areas. Local model explanations explain the sensitivity of individual output predictions to local changes in inputs using gradients (Sundararajan et al., 2017), masks (Petsiuk et al., 2018; Fong et al., 2019), local models (Ribeiro et al., 2016), or Shapley values (Lundberg and Lee, 2017; Chen et al., 2023). Global explanations explain the mechanisms of the overall model at the level of attention patterns (Clark et al., 2019), representations (Belinkov, 2022; Zou et al., 2023a), circuits (Elhage et al., 2021; Wang et al., 2022), or neurons (Bau et al., 2018, 2017; Geva et al., 2023). Models can also be designed to be inherently interpretable (Rudin et al., 2022). Lastly, datasets can be explained using natural language explanations (Narayan et al., 2022), data visualizations (Dibia, 2023), or by training inherently interpretable models (Ustun and Rudin, 2016; Singh et al., 2023).
Model Versions. Models are valuable assets that can be adapted and reused to create new versions. The original or base version is typically a foundation model — a pre-trained model that learns general features from its training dataset, denoted as . By making further adjustments to the training algorithm (), architecture (), or dataset (), subsequent versions of the base model can be developed. Common -based modifications include fine-tuning, parameter-efficient tuning, preference tuning, model stitching, and model editing. Model stitching, for example, involves altering by combining the architectures of two or more models to create a hybrid model (Lenc and Vedaldi, 2015). Model editing (Gandikota et al., 2024; Orgad et al., 2023; Meng et al., 2023, 2022; Mitchell et al., 2022; Sinitsin et al., 2020) focuses on updating certain facts (e.g., changing the name of the current President of a country), making localized adjustments without retraining the entire model. Fine-tuning involves further training a model’s parameters,, to improve performance on specific task(s) or domain(s). For instance, T5 (Raffel et al., 2020) is a pre-trained model that has been trained on a collection of large collection of text (about 750 GB) to perform well on a diverse set of tasks. This has been further fine-tuned with the MIMIC-III (Johnson et al., 2016) and IV (Johnson et al., 2023) dataset to form Clinical-T5 (Lu et al., 2022) to perform better for medical domain-related tasks. In addition to the traditional fine-tuning strategy, parameter-efficient fine-tuning methods have emerged to reduce computational overhead by freezing most of the model’s parameters, only updating a small subset necessary for fine-tuning. For instance, Low Rank Adaption (LoRA) (Hu et al., 2022) is a parameter-efficient fine-tuning method to adapt various models by only updating a low-rank subset (Hu et al., 2022). Preference tuning is another advanced technique, as seen in ChatGPT (OpenAI, [n.d.]) and InstructGPT (Ouyang et al., 2022), which integrates human feedback into the fine-tuning process (Winata et al., 2024). Lastly, due to the emergent ability of models to perform tasks without training (Dong et al., 2022; Liu et al., 2023), newer models leverage prompting as a way to control content generation without needing further updates to their parameters.
Information about model versioning can be inferred even if the model history is unavailable. For example, Mu et al. (Mu et al., 2023) propose a data- and model-driven method to encode “Model DNA" for identifying if a model is a pre-trained version of another, assuming both share the same architecture and training data. However, more challenging cases arise when the target model has a different architecture or is trained on only a subset of the source model’s parameters. Hugging Face recently introduced new metadata fields in their model cards, enabling users to specify the base model and explain how it has been modified. This metadata generates a model tree, linking related models by their extensions. However, its accuracy depends on reliable documentation, and older models lack this data. Research on reconstructing and verifying relationship is emerging (Horwitz et al., 2024; Sun et al., 2025; Zhu et al., 2025), such as Horwitz et al. (2024)’s approach using weight similarities, though this approach is limited to known models with a single base version.
Privacy and Safety. Models are vulnerable to the disclosure of private information (Shokri et al., 2017; Carlini et al., 2021) and adversarial results (Zou et al., 2023b; Li et al., 2025) when attacked. Hence, initiatives such as Privacy Preserving Machine Learning (Ruehle et al., 2021) exist to understand, measure, and mitigate such risks. As a result, techniques such as differential privacy (Dwork et al., 2006; Yan et al., 2024), data sanitization (Staab et al., 2025; Dou et al., 2024) and robust prompt optimization (Zhou et al., 2024) have been utilized to defend against such attacks. These methods generally aim to obscure or eliminate private information while detecting and preventing attempts to jailbreak or manipulate the model and its output. However, this can create a false sense of privacy as defense techniques can continue to be compromised with other attack schemes (Xin et al., 2024).
Beyond privacy, ensuring AI systems are fundamentally safe is another critical challenge. Community-driven efforts focus on building safe AI (Ortega et al., 2018), aiming to specify, verify, and ensure that the models behave as intended. A recent approach leverages neuroscience concepts, such as representation engineering, to enhance AI transparency (Andy Zou, 2023; Mineault et al., 2024) and improve our understanding of traits like honesty, power seeking tendencies, and morality in models. However, this direction is still in its early stages, with open challenges including determining the range of representations of concepts and functions that can be extracted from models and developing scalable evaluation methods to better inform users about model safety.
Benchmarking. Benchmarking plays a crucial role in evaluating model performance for specific tasks and remains a well-established research area, however benchmarks for newer topics such as model attribution and versioning are urgently needed. Developers frequently report model performance using standardized benchmarks, making it a routine part of model assessment. Broadly, there are two primary types of benchmarking: (1) evaluating how well a model, , approximates a ground truth distribution, and (2) assessing a model’s performance against a targeted evaluation metric. In classification tasks, models are typically evaluated using accuracy, as ground-truth labels provide a direct assessment of the model’s output. Additionally, confusion matrices offer insights into the types of errors made. For text generation tasks, perplexity is a widely used metric, and several popular benchmarks exist to assess performance (Hendrycks et al., 2021a, b; Rein et al., 2024). In the case of image generation, the Fréchet Inception Distance (FID) (Heusel et al., 2017) is a common metric and COCO (Lin et al., 2014) and VQAv2 (Goyal et al., 2017) are some notable benchmarks in this domain. Beyond traditional performance metrics, benchmarking also considers biases related to protected attributes (Schramowski et al., 2023), as well as the environmental impact (Lacoste et al., 2019) of model training and deployment. Model lake benchmarks lack large-scale, publicly available datasets that mimic realistic conditions of diverse models in model lakes. There has been preliminary efforts by Lu et al. (2023), where they released a benchmark for model search with 259 publicly available image generative models and 1000 customized text-to-image diffusion models. Model Zoo (Schürholt et al., 2022) is another relatively large-scale dataset, but it is limited to vision models. Creating large-scale model lake benchmark that integrates various modalities (such as text, images, and audio) remains a research challenge.
5. Research Roadmap
Building on current research and identified gaps outlined in Section 4, this section presents a vision to address these challenges. A model lake is illustrated in Figure 2, where rather than APIs, we envision data scientists interaction with the lake using queries.
Benchmarking. As discussed in Section 4, benchmarking, while a well-established research area, is under-explored in model lakes. An important topic is the development of lifelong benchmarks (Prabhu et al., 2024) that can address increasingly complex and novel scenarios as models continue to evolve in capability and diversity. In addition, there is a need to develop benchmarks specific to model lake tasks. For instance, to advance research in model attribution, a comprehensive benchmark dataset is needed—one that includes labeled model parameters, architectures, and detailed transformation records (e.g., fine-tuning, model editing). This benchmark can be extended to model versioning by adding data that tracks the previous and subsequent versions of models.
Model Inference. Deploying effective solutions for model lake tasks is a challenge with respect to usability and scalability. To improve usability, the model inference component involves identifying appropriate benchmarks and generating relevant prompts, as well as selecting suitable models (target models or meta-models) to address a user query. While users can manually run prompts and select models, this approach is prone to errors and suboptimal outcomes, especially if users lack the expertise to use models effectively. For example, a classifier’s behavior may be misinterpreted if a user does not understand the type of data is was trained or the input it expects. To mitigate this, the proposed model lake tasks can incorporate additional perspectives, such as intrinsic model properties (e.g., weights), to provide insights and guide users towards a more accurate interpretations and applications. This search and generation process can also be automated using an AI agent. By applying benchmarks and model(s) in question, we combine research efforts of benchmarking and attribution questions, which can explain a model’s behavior and output.
Indexer. A central component of a model lake is the indexer, which would be used to embed and provide scalable sublinear search over the model embeddings. The indexer can use different ranking functions tailored to the task. Wang et al. (2024b); Achille et al. (2019) have introduced an approach for generating model and task embeddings. However, search methods must scale to handle millions of models and adapt to newer models with advanced capabilities (Wei et al., 2022). Indices like HNSW (Hierarchical New Small World) (Malkov and Yashunin, 2020), have proven effective in practice in indexing high-dimensional embeddings enabling fast nearest-neighbor search (including over data lakes (Fan et al., 2023b). However, HNSW provides no formal guarantees on correctness and its use in model lakes remains under-explored. Effective embedding of models is crucial for accurate comparison and ranking by the indexer. Wang et al. (2024b) explores this but there work is not inclusive of all model types. A robust system should support diverse embeddings to ensure indexing effectiveness. Many of the model lake tasks will benefit from hybrid approach, that indexes both metadata and model embeddings – for example, related model search can combine well-chosen model embeddings representing important intrinsic model features with search over verified models cards. Similarly, in versioning, embeddings and their associated rankings can aid in identifying parent-child relationships and assessing the distinctiveness of each model relative to others.
Weight-Space Modeling. Weight-Space modeling is a hyper-representation learning approach where a neural network is trained to process weights of other models (Eilertsen et al., 2020). This method can be useful for making distinctions between models, especially in complex scenarios like model stitching, where similar models with multiple shared parent models need to be distinguished. This is a promising direction. Zhou et al. (2025), for instance, reveals a linear connection between fine-tuned models. This approach could also facilitate dynamic selection of benchmarks for performance measurement by learning from previous relationships between datasets and models. The primary challenge in weight-space modeling is identifying the most relevant aspects of the model for training another model to uncover patterns, while simultaneously considering the weight-space model architecture (Navon et al., 2023). It is also crucial to ensure sufficient data diversity to avoid overfitting or underfitting.
Interpretability. Interpretability methods can be most useful for the model attribution task where users must understand the origin of model behavior, for example, when detecting knowledge changes (Youssef et al., 2024), or when unlearning learned knowledge (Jang et al., 2023; Eldan and Russinovich, 2023; Gandikota et al., 2023; Belrose et al., 2023; Kumari et al., 2023). Attribution questions also apply to the source model’s architecture. Approaches like circuit discovery (Wang et al., 2022; Conmy et al., 2023; Elhage et al., 2021) can help identify the computational origin of model behavior. Model inversion can be used to recover an input prompt given an output (Morris et al., 2023, 2024; Huang et al., 2024). These methods are also a promising route for understanding the impact of training data on internal model states. In recent work, Sharkey et al. (2025) present a list of open questions in interpretability, many of which are relevant to the problem of model lakes.
Holistic Management of Models and Data. Effective model lakes need to integrate advances in data lakes in a holistic way given the reliance of models on data as their fuel. Many model management tasks include (in part) an analysis of training or input data. Hence, traditional data lake concerns such as data provenance, data versioning, and related issues still apply. Thus, in addition to new tasks specific to model lakes, we must also account for data lake tasks when dealing with the data used by or generated from these models. And integrating these tasks will be important. As a simple example, when searching for models trained on a dataset, users may want to find models trained on versions of the dataset.
6. MODEL LAKE TASKS: APPLICATIONS
Documentation Generation. To streamline the creation of detailed model cards, engineers can leverage model lake tasks to generate a rough draft of the required documentation. This application is similar to discovering and annotating metadata in data lakes (Akhtar et al., 2024; Sawadogo and Darmont, 2021; Zhao et al., 2021; Korini and Bizer, 2023; Fan et al., 2023a). Here’s how the process might work: upon uploading a model to the model lake, state-of-the-art techniques for tasks like attribution, versioning, benchmarking, and others can automatically analyze and map the model’s relationships to other models in the model lakes. The engineer can review and either accept or modify these generated mappings, especially if any information appears inaccurate, creating an initial version of the model details section. Additionally, the engineer can assess the model’s robustness by testing its performance against relevant benchmarks, including metrics like task accuracy, environmental impact, and more. Based on these test results, key sections of the model card, such as intended use and performance metrics, can be auto-populated for a faster, more reliable drafting process. It can alert developers to model risks, as shown by Wang et al. (2024a), who demonstrate how model versioning helps warn downstream model users of upstream model risks.
Auditing. Policymakers have recently proposed AI safety and compliance regulations (House, 2022; Tabassi, 2023, 2024) aimed at fostering more responsible and accountable AI models (Casper et al., 2024). The model document generation application procedure can be repurposed for auditing by creating a template questionnaire (Golpayegani et al., 2024) and using the information from the model lake to generate a draft response with proof or explanation about how a requirement is fulfilled. This process can incorporate privacy-related technical solutions, where insights from model lake tasks help identify vulnerabilities across related models and their successor versions. It can also aid in attributing sensitive data that the model may have access to, highlighting potential exposure risks.
Data and Model Citation. Data citation helps stakeholders identify the source, ownership, and authorship of the data used for a particular analysis. It is important to use proper data citation because the structure and contents of the database can evolve (Buneman et al., 2016). Hence, this problem is also extended to machine learning, since a large part of model training is its dataset. Thus, the task of citing data for data lakes remains relevant for documenting the training data used in the creation of the model. Similarly, model citation is essential, as users can further train the model or use its outputs for consumption or additional training. One proposed solution to identify generated output is the use of watermarks (Kirchenbauer et al., 2023). In addition, model versioning tasks provide crucial documentation, allowing researchers, engineers, and developers to refer to the exact version of the model used for training or content generation. If a particular model is used, the platform would refer to its versioning graph and generate a citation with the model version and timestamp of the graph. Upon any updates of the graph, a new citation would be generated with the updated version and timestamp. This would be useful for also citing any generated content from this model.
Model Search. In Example 1.1, we illustrate a potential scenario where a user seeks the most suitable model for summarizing and simplifying legal documents. As the model search task within the model lake evolves, we aim for users to be able to write declarative queries and retrieve a set of models ranked by their suitability for the specified task. Query example include "Find all models trained on this corpus of US Supreme Court cases" or "Find models that out perform Model X on Benchmark Y". Given a search task, the model lake framework can map the task function to a suitable indexer and run that indexer to retrieve top-ranked models. Whether the user is a technical or non-technical consumer, researcher, or engineer, they would be able to click on a model to view its model card. Attribution reveals how training data or learned concepts influence outputs, while version management clarifies the model’s training process, lineage, and differences, enhancing transparency. Benchmarking evaluates the model’s robustness on related tasks, addressing completeness. Together, these allow users to make informed model choices.